# Unified Smooth Vector Graphics: Modeling Gradient Meshes and Curve-based Approaches Jointly as Poisson Problem

### v1

@@ 1 Introduction
The field of smooth vector graphics describes image content based on geometric primitives, such as curves or meshes. This is in contrast to raster graphics, which store the color of each pixel explicitly. Smooth vector graphics are a popular tool for the modeling of scale-independent content, such as icons, schematic illustrations, or web content (Hsiao et al., 2023), since images can be synthesized on any output resolution. Basic vector graphics primitives are standardized in the SVG format, which includes the most basic shapes and color gradients. In research, two conceptually different approaches have been developed to extend the artistic control in smoothly-shaded regions. On the one hand, there are gradient meshes (Sun et al., 2007; Lai et al., 2009), which interpolate colors from tensor product surfaces efficiently. And on the other hand, there are curve-based approaches such as diffusion curves (Orzan et al., 2008; Jeschke, 2016) and Poisson curves (Hou et al., 2018) that model the final image as solution to a diffusion problem from user-defined boundary conditions, e.g., curve colors. Both approaches remained incompatible, and research on editing, rasterization, and vectorization, has continued separately on the two research threads (Tian and Günther, 2023). With this paper, we propose a mathematical formulation that allows combining gradient meshes and curve-based approaches consistently for the first time. Fig. 1 gives an example, where gradient meshes are combined with diffusion curves and Poisson curves. To maximize the compatibility with existing methods we take a set of gradient meshes, diffusion curves, and Poisson curves as input, which are well-known primitives that the user interacts with on the front-end. To unify the modeling, we convert those representations first into an intermediate patch representation that specifies for each patch a target Laplacian function and boundary conditions. This intermediate representation is not visible on the user interface and is only used internally to assemble the Poisson problems. We propose to model the pixel color within gradient meshes as solution to a partial differential equation (PDE) in order to match their mathematical treatment with that of curve-based methods. In the conversion process to our patch representation, geometric intersections of gradient meshes and curves get resolved automatically to yield a non-overlapping domain decomposition. In our new formulation, additional types of boundary conditions are added to the existing primitives, which extends the artistic freedom. For example, Neumann conditions can be specified on diffusion curves, as in Bang et al. (2023). To synthesize a raster image for a given output resolution, we rasterize boundary conditions and Laplacians for the respective patches and compute the final image as solution to a Poisson problem. We evaluate the method on various test scenes containing gradient meshes and curve-based primitives. In summary, our contributions are:
@@
For the first time, our approach enables the unified mathematical treatment of gradient meshes and curve-based methods, which have been developed independently for over a decade. Since we take existing gradient mesh and curve formulations as input, our approach achieves high compatibility with existing content creation pipelines. Further, the final algorithm leads to a standard Poisson problem, for which highly optimized numerical solvers readily exist. The benefits of combining gradient meshes and curve-based formulations in the same scene are shown in Fig. 2. While curve-based approaches are useful to create intricate details, it is easier to create larger color gradients with mesh-based approaches (Tian and Günther, 2023). By combining both methods, we are able to generate expressive results that are difficult to obtain with either representation on its own.
@@ 2 Background
In this section, we formally introduce the notation used to describe mesh-based and curve-based vector graphics primitives. For a detailed coverage, we refer to a recent survey (Tian and Günther, 2023).
@@ 2.2 Mesh-based Methods
To simplify the modeling, the mixed partial derivatives are commonly set to zero, i.e., $\partial_{u}\partial_{v}f_{i,j}=0$, and thus, a function $f(u,v)$ can be controlled by the user by specifying the values $f_{i,j}$ and the tangent handles $\partial_{u}f_{i,j}$ and $\partial_{v}f_{i,j}$ at the four corners $i,j\in\{0,3\}$, see Fig. 3a. A gradient mesh consists of multiple Ferguson patches that are stitched together with suitable continuity condition. The functions $f(u,v)$ that are modeled by the gradient mesh are the position $\mathbf{x}(u,v)$ and the color $\mathbf{c}(u,v)$. Thus, the degrees of freedom are $\mathbf{x}_{i,j}$, $\mathbf{c}_{i,j}$ and the tangent handles $\partial_{u}\mathbf{x}_{i,j}$, $\partial_{u}\mathbf{c}_{i,j}$ and $\partial_{v}\mathbf{x}_{i,j}$, $\partial_{v}\mathbf{c}_{i,j}$ at the four corners $(i,j)\in\{(0,0),(0,3),(3,0),(3,3)\}$. To model holes in surfaces, Lai et al. (2009) proposed to split gradient meshes along discontinuities. Barendrecht et al. (2018) enabled local refinement of cubic rectangular patches. In line with many other papers (Xiao et al., 2012, 2015; Wan et al., 2018; Wei et al., 2019), we likewise base our description of gradient meshes on Sun et al. (2007), see Eq. (4).
@@
As an alternative to rectangular meshes, triangle-based patches have been explored, where the patch boundaries are curved edges (Xia et al., 2009; Xiao et al., 2022). Control over the continuity across adjacent patches is desirable (Liao et al., 2012; Zhou et al., 2014). To this end, quadratic triangle configuration B-splines (TCBs) (Cao et al., 2019; Schmitt, 2019) have been used by Zhu et al. (2022) to model color variations while staying $C^{1}$ continuous. An example for irregular patches are the polygonal patches of Swaminarayan and Prasad (2006), who used constant colors in the interior. With the introduction of Bézigons (Yang et al., 2015) closed regions have been bounded by Bézier curves. For color interpolation in the interior, several options are available (Hettinga et al., 2019), including cubic mean value coordinates (Li et al., 2013). Since the methods above use differentiable interpolants, they are compatible with our approach.
@@ 2.3 Curve-based Methods
Follow-up research on rendering (Pang et al., 2011; Sun et al., 2014; Bang et al., 2023), vectorization (Zhao et al., 2017; Lu et al., 2019) and editing (Jeschke et al., 2011; Lu et al., 2020) utilized this formulation. To provide more control over the color gradients, Bezerra et al. (2010) controlled the strength and direction of the diffusion. Alternatively, Sun et al. (2012) introduced textures in diffusion curves. Several renderers have been proposed by now that are able to rasterize diffusion curves, including ray tracing approaches (Bowers et al., 2011; Prévost et al., 2015), Monte Carlo methods (Sawhney and Crane, 2020; Sawhney et al., 2023; Sugimoto et al., 2023), and boundary element methods (Ilbery et al., 2013; Bang et al., 2023; Chen et al., 2024).
@@
Thus, rather than looking for the smoothest possible color gradient (via $\mathbf{f}(\mathbf{x})=0$), bumps and discontinuities can be introduced in the color gradient, which is locally controlled by the user by drawing curve and region primitives, see Fig. 3c. Non-constant Laplacians in Poisson regions have been described by raster images, and were utilized for color transfer between images (Fu et al., 2019) and for image vectorization (Fu et al., 2024). Alternatively, Finch et al. (2011) proposed to constrain high-order derivatives, seeking solutions that are harmonic in their Laplacian domain, i.e., $\Delta^{2}\mathbf{c}(\mathbf{x})=\mathbf{0}$. With this formulation, extrapolation may create new extrema at unconstrained positions (Finch et al., 2011; Boyé et al., 2012), which requires manual adjustment of control points (Finch et al., 2011) or a non-linear optimization (Jacobson et al., 2012). Further, the diffused color range is tied to the scene geometry, which may lead to oversaturation (Jeschke, 2016) when scaling objects. In our work, we concentrate on the Poisson formulation in Eq. (6), which includes the diffusion curves in Eq. (5) as special case. Poisson vector graphics (Hou et al., 2018) have led to a new line of work on the control over color gradients within smooth vector graphics. In this paper, we seek to utilize the well-established gradient meshes (Sun et al., 2007), which so far have not been incorporated into this framework.
(a) Input Primitives
(b) Edge Graph
(c) Unified Patches
(d) Image Synthesis
@@ 3 Method Overview
    Image Synthesis. Given the patches with their boundary conditions, we solve the Poisson problem for each patch independently with an off-the-shelf Poisson solver.
In the following, the four steps are explained in more detail.
@@ 4 Input Primitives
Input to our approach is a collection of extended gradient meshes $\mathcal{G}$, a collection of extended diffusion curves $\mathcal{D}$ and a collection of Poisson curves $\mathcal{P}$. Those three types of input primitives are drawn by the user and can be edited and controlled by manipulation of control points, tangent handles, and color stops, identical to the editing processes of conventional diffusion curves, gradient meshes, and Poisson curves (Tian and Günther, 2023). In the following, we explain how the primitives are defined and how they are extended in our framework.
@@ 5.2 Construction
We begin with an empty graph, i.e., $\mathcal{V}=\varnothing$, $\mathcal{E}=\varnothing$. For each input boundary curve $(\mathbf{x}(t),\mathcal{B}_{l},\mathcal{B}_{r})\in\mathcal{G}\cup\mathcal{D}$ from all extended gradient meshes $\mathcal{G}$ and extended diffusion curves $\mathcal{D}$, we take the following steps:
@@
    If there is any other existing point $\mathbf{v}_{k}\in\mathcal{V}$ with $\|\mathbf{v}_{i}-\mathbf{v}_{k}\|^{2}<\epsilon$ or $\|\mathbf{v}_{j}-\mathbf{v}_{k}\|^{2}<\epsilon$, then merge nearby points.
@@
An example is given in Fig. 8. An inserted input boundary curve may have multiple intersections with other curves or with itself. The intersections are resolved by inserting vertices until no further intersections can be found. With this, the graph is brought back to a consistent state and the next input boundary curve can be processed.
@@ 5.3 Implementation Detail
During the graph construction process, input boundary curve intersections need to be calculated. To accelerate the intersection tests, we internally discretize the input boundary curves into piece-wise linear polylines in a bottom-up construction using the Ramer-Douglas-Peucker algorithm (Ramer, 1972; Douglas and Peucker, 1973). Finding curve intersections then reduces to a line-line intersection test, which we accelerate by forming a bounding volume hierarchy. When a new curve is added by the user or when the control points of an existing curve are moved, the curve is re-discretized and the bounding volume hierarchy is refit.
@@ 6.2 Construction
The valence $|\mathbf{v}_{i}|$ of vertex $\mathbf{v}_{i}\in\mathcal{V}$ in the graph, i.e., the number of edges connected to it, tells how often the vertex can be visited when traversing the graph in the search for closed patches, namely $2\cdot|\mathbf{v}_{i}|$ (with the exception of image boundary vertices for which it is $2\cdot|\mathbf{v}_{i}|-2$). Starting from the first vertex and its first edge in the graph, we traverse the edges from vertex to vertex tracing out a sequence of patch boundary curves that encloses a patch. A closed patch is formed by always turning right when visiting a junction, i.e., a vertex. This is illustrated in Fig. 10 (left). For the required angle measurements, the curve tangents $\dot{\mathbf{x}}(t)$ are evaluated at the end vertex of the edge, i.e., at $t_{0}$ or $t_{1}$, respectively. The edge traversal terminates when a loop was found, i.e., when the traversal returns to the start vertex. By recording how often a vertex was visited, we repeat the process until all patches are inserted and no vertex is expecting another visit.
@@
The patch Laplacian $\mathbf{f}(\mathbf{x})$ defines the color gradient inside a patch, which is controlled by evaluating and adding the Laplacian $\mathbf{f}_{\mathbf{p}(t)}(\mathbf{x})$ of Poisson curves $\mathbf{p}(t)\in\mathcal{P}$ and by sampling the color derivatives of gradient meshes, cf. Section $2.$
[…]
If one (and only one) patch boundary curve of a patch is part of a gradient mesh interior, we add the Laplacian of the gradient mesh $\Delta_{\mathbf{x}}\mathbf{c}(\mathbf{u}(\mathbf{x}))$, which can be evaluated analytically as explained in Appendix A (including the coordinate transformation $\mathbf{u}(\mathbf{x})$ that takes image coordinates to a $(u,v)$ coordinate in the gradient mesh). When a patch is formed from a gradient mesh, then the diffused color gradient inside the patch matches exactly the color gradient of the interpolated gradient mesh. If two gradient meshes overlap, then there is a region in which two Laplacian functions are defined. In this case, it is up to the user to decide, which Laplacian to use, for example a linear combination of both, or the one of the gradient mesh that is ’on top’, i.e., which was inserted last. To keep the implementation simple, we decided to use neither of them. Thus, if (a part of) a patch is not bound by exactly one gradient mesh, then the Laplacian of the function is only the sum of Poisson curves’s Laplacians. It there are no Poisson curves, then the patch Laplacian is homogeneous $\mathbf{f}(\mathbf{x})=\mathbf{0}$, which is equivalent to the setting for regular diffusion curves, cf. Eq. (5).

### v2

@@ 1 Introduction
The field of smooth vector graphics describes image content based on geometric primitives, such as curves or meshes. This is in contrast to raster graphics, which store the color of each pixel explicitly. Smooth vector graphics are a popular tool for the modeling of scale-independent content, such as icons, schematic illustrations, or web content (Hsiao et al., 2023), since images can be synthesized on any output resolution. Basic vector graphics primitives are standardized in the SVG format, which includes the most basic shapes and color gradients. In research, two conceptually different approaches have been developed to extend the artistic control in smoothly-shaded regions. On the one hand, there are gradient meshes (Sun et al., 2007; Lai et al., 2009), which interpolate colors from tensor product surfaces efficiently. And on the other hand, there are curve-based approaches such as diffusion curves (Orzan et al., 2008; Jeschke, 2016) and Poisson curves (Hou et al., 2018) that model the final image as solution to a diffusion problem from user-defined boundary conditions, e.g., curve colors. Both approaches remained incompatible, and research on editing, rasterization, and vectorization, has continued separately on the two research threads (Tian and Günther, 2024). With this paper, we propose a mathematical formulation that allows combining gradient meshes and curve-based approaches consistently for the first time. Fig. 1 gives an example, where gradient meshes are combined with diffusion curves and Poisson curves. To maximize the compatibility with existing methods we take a set of gradient meshes, diffusion curves, and Poisson curves as input, which are well-known primitives that the user interacts with on the front-end. To unify the modeling, we convert those representations first into an intermediate patch representation that specifies for each patch a target Laplacian function and boundary conditions. This intermediate representation is not visible on the user interface and is only used internally to assemble the Poisson problems. We propose to model the pixel color within gradient meshes as solution to a partial differential equation (PDE) in order to match their mathematical treatment with that of curve-based methods. In the conversion process to our patch representation, geometric intersections of gradient meshes and curves get resolved automatically to yield a non-overlapping domain decomposition. In our new formulation, additional types of boundary conditions are added to the existing primitives, which extends the artistic freedom. For example, Neumann conditions can be specified on diffusion curves, as in Bang et al. (2023). To synthesize a raster image for a given output resolution, we rasterize boundary conditions and Laplacians for the respective patches and compute the final image as solution to a Poisson problem. We evaluate the method on various test scenes containing gradient meshes and curve-based primitives. In summary, our contributions are:
@@
  • We provide an open source implementation that reads the scenes, constructs the patches, and renders them.
  • We compare the unified scenes with vectorizations using gradient meshes only, and using diffusion curves only.
For the first time, our approach enables the unified mathematical treatment of gradient meshes and curve-based methods, which have been developed independently for over a decade. Since we take existing gradient mesh and curve formulations as input, our approach achieves high compatibility with existing content creation pipelines. Further, the final algorithm leads to a standard Poisson problem, for which highly optimized numerical solvers readily exist. The benefits of combining gradient meshes and curve-based formulations in the same scene are shown in Fig. 2. While curve-based approaches are useful to create intricate details, it is easier to create larger color gradients with mesh-based approaches (Tian and Günther, 2024). By combining both methods, we are able to generate expressive results that are difficult to obtain with either representation on its own.
@@ 2 Background
In this section, we formally introduce the notation used to describe mesh-based and curve-based vector graphics primitives. For a detailed coverage, we refer to a recent survey (Tian and Günther, 2024).
@@ 2.2 Mesh-based Methods
To simplify the modeling, the mixed partial derivatives are commonly set to zero, i.e., $\partial_{u}\partial_{v}f_{i,j}=0$, and thus, a function $f(u,v)$ can be controlled by the user by specifying the values $f_{i,j}$ and the tangent handles $\partial_{u}f_{i,j}$ and $\partial_{v}f_{i,j}$ at the four corners $i,j\in\{0,3\}$, see Fig. 3a. A gradient mesh consists of multiple Ferguson patches that are stitched together with suitable continuity condition. The functions $f(u,v)$ that are modeled by the gradient mesh are the position $\mathbf{x}(u,v)$ and the color $\mathbf{c}(u,v)$. Thus, the degrees of freedom are $\mathbf{x}_{i,j}$, $\mathbf{c}_{i,j}$ and the tangent handles $\partial_{u}\mathbf{x}_{i,j}$, $\partial_{u}\mathbf{c}_{i,j}$ and $\partial_{v}\mathbf{x}_{i,j}$, $\partial_{v}\mathbf{c}_{i,j}$ at the four corners $(i,j)\in\{(0,0),(0,3),(3,0),(3,3)\}$. To model holes in surfaces, Lai et al. (2009) proposed to split gradient meshes along discontinuities. Barendrecht et al. (2018) enabled local refinement of cubic rectangular patches. Baksteen et al. (2021) fitted mesh colors (Yuksel et al., 2010) for colorization of gradient meshes. In line with many other papers (Xiao et al., 2012, 2015; Wan et al., 2018; Wei et al., 2019), we likewise base our description of gradient meshes on Sun et al. (2007), see Eq. (4).
@@
As an alternative to rectangular meshes, triangle-based patches have been explored, where the patch boundaries are straight (Lawonn and Günther, 2019) or curved edges (Xia et al., 2009; Xiao et al., 2022; Wang et al., 2024). Control over the continuity across adjacent patches is desirable (Liao et al., 2012; Zhou et al., 2014). To this end, quadratic triangle configuration B-splines (TCBs) (Cao et al., 2019; Schmitt, 2019) have been used by Zhu et al. (2022) to model color variations while staying $C^{1}$ continuous. An example for irregular patches are the polygonal patches of Swaminarayan and Prasad (2006), who used constant colors in the interior. With the introduction of Bézigons (Yang et al., 2015) closed regions have been bounded by Bézier curves. For color interpolation in the interior, several options are available (Hettinga et al., 2019, 2021), including cubic mean value coordinates (Li et al., 2013). Since the methods above use differentiable interpolants, they are compatible with our approach.
@@ 2.3 Curve-based Methods
Follow-up research on rendering (Pang et al., 2011; Sun et al., 2014; Bang et al., 2023), vectorization (Zhao et al., 2017; Lu et al., 2019) and editing (Jeschke et al., 2011; Lu et al., 2020) utilized this formulation. To provide more control over the color gradients, Bezerra et al. (2010) controlled the strength and direction of the diffusion. Alternatively, Sun et al. (2012) introduced textures in diffusion curves. Several renderers have been proposed by now that are able to rasterize diffusion curves, including ray tracing approaches (Bowers et al., 2011; Prévost et al., 2015), Monte Carlo methods (Sawhney and Crane, 2020; Sawhney et al., 2023; Sugimoto et al., 2023; Li et al., 2023; Miller et al., 2023; Sugimoto et al., 2024; Miller et al., 2024b; Czekanski et al., 2024), and boundary element methods (Ilbery et al., 2013; Bang et al., 2023; Chen et al., 2024). Lately, Monte Carlo methods have been phrased in a differentiable manner, which is useful for inverse problems such as vectorizations (Yilmazer et al., 2024; Yu et al., 2024; Miller et al., 2024a).
@@
Thus, rather than looking for the smoothest possible color gradient (via $\mathbf{f}(\mathbf{x})=0$), bumps and discontinuities can be introduced in the color gradient, which are controlled locally by the user by drawing curve and region primitives, see Fig. 3c. Non-constant Laplacians in Poisson regions have been described by raster images, and were utilized for color transfer between images (Fu et al., 2019) and for image vectorization (Fu et al., 2024). Alternatively, Finch et al. (2011) proposed to constrain high-order derivatives, seeking solutions that are harmonic in their Laplacian domain, i.e., $\Delta^{2}\mathbf{c}(\mathbf{x})=\mathbf{0}$. With this formulation, extrapolation may create new extrema at unconstrained positions (Finch et al., 2011; Boyé et al., 2012), which requires manual adjustment of control points (Finch et al., 2011) or a non-linear optimization (Jacobson et al., 2012). Further, the diffused color range is tied to the scene geometry, which may lead to oversaturation (Jeschke, 2016) when scaling objects. In our work, we concentrate on the Poisson formulation in Eq. (6), which includes the diffusion curves in Eq. (5) as special case. Poisson vector graphics (Hou et al., 2018) have led to a new line of work on the control over color gradients within smooth vector graphics. In this paper, we seek to utilize the well-established gradient meshes (Sun et al., 2007), which so far have not been incorporated into this framework.
@@ 3 Method Overview
    Image Synthesis. Given the patches with their boundary conditions, we solve the Poisson problem for each patch with an off-the-shelf Poisson solver.
A C++ reference implementation of our method can be found at (Tian and Günther, 2025). In the following, the four steps are explained in more detail.
(a) Input Primitives
(b) Edge Graph
(c) Unified Patches
(d) Image Synthesis
@@ 4 Input Primitives
Input to our approach is a collection of extended gradient meshes $\mathcal{G}$, a collection of extended diffusion curves $\mathcal{D}$ and a collection of Poisson curves $\mathcal{P}$. Those three types of input primitives are drawn by the user and can be edited and controlled by manipulation of control points, tangent handles, and color stops, identical to the editing processes of conventional diffusion curves, gradient meshes, and Poisson curves (Tian and Günther, 2024). In the following, we explain how the primitives are defined and how they are extended in our framework.
@@ 5.2 Construction
An example of the edge graph construction is given in Fig. 8. We begin with an empty graph, i.e., $\mathcal{V}=\varnothing$, $\mathcal{E}=\varnothing$. For each input boundary curve $(\mathbf{x}(t),\mathcal{B}_{l},\mathcal{B}_{r})\in\mathcal{G}\cup\mathcal{D}$ from all extended gradient meshes $\mathcal{G}$ and extended diffusion curves $\mathcal{D}$, we take the following steps:
@@
    If there is any other existing point $\mathbf{v}_{k}\in\mathcal{V}$ with $\|\mathbf{v}_{i}-\mathbf{v}_{k}\|^{2}<\tau$ or $\|\mathbf{v}_{j}-\mathbf{v}_{k}\|^{2}<\tau$, then merge nearby points.
@@
An inserted input boundary curve may have multiple intersections with other curves or with itself. The intersections are resolved by inserting vertices until no further intersections can be found. With this, the graph is brought back to a consistent state and the next input boundary curve can be processed. The parameter $\tau$ allows for closing gaps that would otherwise cause unwanted color leaking. Fig. 9 illustrates the snapping of an edge endpoint onto an existing curve. For this, the closest point on the existing edge is determined, which is either one of the end points or a point on the edge. In case of the latter, the existing edge is subdivided, and the vertex of the new edge is snapped onto the curve. Fig. 10 shows renderings for different thresholds $\tau$ for a simple test scene.
input configuration
find closest point
snap $\mathbf{v}_{4}$ onto curve
Figure 9: When the endpoint of the next edge ($\mathbf{e}_{34}$, green) is close to an already existing edge ($\mathbf{e}_{12}$, black), then we snap the endpoint of the new edge onto the existing edge.
$\tau=0.0$
$\tau=0.1$
$\tau=0.2$
Figure 10: Here, the snapping threshold $\tau$ is varied. On the left, the result without clean-up is shown. In the middle, the two endpoints at the bottom merged. On the right, the gap at the top was closed. This simple test scene is defined over the domain $0.5,2.5]^{2}$.
@@ 5.3 Implementation Detail
During the graph construction process, input boundary curve intersections need to be calculated. To accelerate the intersection tests, we internally discretize the input boundary curves into piece-wise linear polylines in a top-down construction that is based on the Ramer-Douglas-Peucker algorithm (Ramer, 1972; Douglas and Peucker, 1973). However, instead of starting from a polyline, we start directly with the cubic Bézier curve. In Fig. 11(a), we define a line that connects the first and last vertex of the curve. In (b), we determine the curve point, which has the largest vertical distance from the line. This is done recursively using Bézier clipping (Sederberg and Nishita, 1990). In (c), we insert a new vertex at that point and recursively proceed on the segments left and right. In (d), we see the result after termination, which happens when the distance of the farthest point falls below an error threshold $\epsilon$. In (e), we can see that it is not sufficient to only measure the vertical distance to the line (blue). Instead, we also compute the (signed) horizontal distance from the start and end point (green). For subdivision, the maximum of the three distances is computed. In (g), the result is shown, which successfully approximates the curve. Examples of different discretizations are shown in Fig. 12 for a simple test scene. Finding curve intersections then reduces to a line-line intersection test. When a new curve is added by the user or when the control points of an existing curve are moved, the curve is re-discretized.
Figure 11: Illustration of the curve discretization method. In (a)-(d), the recursive subdivision is shown, while (e)-(f) illustrates at an example why measuring vertical distances alone is not sufficient.
$\epsilon=0.01$
$\epsilon=0.1$
$\epsilon=0.5$
Figure 12: Here, the results for different curve discretizations are shown, which are obtained when choosing different error thresholds $\epsilon$. This simple test scene is defined over the domain $0.5,2.5]^{2}$.
@@ 6.2 Construction
The valence $|\mathbf{v}_{i}|$ of vertex $\mathbf{v}_{i}\in\mathcal{V}$ in the graph, i.e., the number of edges connected to it, tells how often the vertex can be visited when traversing the graph in the search for closed patches, namely $2\cdot|\mathbf{v}_{i}|$ (with the exception of image boundary vertices for which it is $2\cdot|\mathbf{v}_{i}|-2$). Starting from the first vertex and its first edge in the graph, we traverse the edges from vertex to vertex tracing out a sequence of patch boundary curves that encloses a patch. A closed patch is formed by always turning right when visiting a junction, i.e., a vertex. This is illustrated in Fig. 14 (left). For the required angle measurements, the curve tangents $\dot{\mathbf{x}}(t)$ are evaluated at the end vertex of the edge, i.e., at $t_{0}$ or $t_{1}$, respectively. The edge traversal terminates when a loop was found, i.e., when the traversal returns to the start vertex. By recording how often a vertex was visited, we repeat the process until all patches are inserted and no vertex is expecting another visit. Fig. 15 shows a number of exemplary edge graphs. In the following, we list the corresponding traversal sequences.
(a) Traverse outer edge of a closed domain: $\mathbf{v}_{1}$ $\rightarrow$ $\mathbf{e}_{12}$ $\rightarrow$ $\mathbf{v}_{2}$ $\rightarrow$ $\mathbf{e}_{23}$ $\rightarrow$ $\mathbf{v}_{3}$ $\rightarrow$ $\mathbf{e}_{34}$ $\rightarrow$ $\mathbf{v}_{4}$ $\rightarrow$ $\mathbf{e}_{41}$ $\rightarrow$ $\mathbf{v}_{1}$
(b) Traverse single edge: $\mathbf{v}_{1}$ $\rightarrow$ $\mathbf{e}_{12}$ $\rightarrow$ $\mathbf{v}_{2}$ $\rightarrow$ $\mathbf{e}_{12}$ $\rightarrow$ $\mathbf{v}_{1}$
(c) Traverse sequence of edges: $\mathbf{v}_{1}$ $\rightarrow$ $\mathbf{e}_{12}$ $\rightarrow$ $\mathbf{v}_{2}$ $\rightarrow$ $\mathbf{e}_{23}$ $\rightarrow$ $\mathbf{v}_{3}$ $\rightarrow$ $\mathbf{e}_{23}$ $\rightarrow$ $\mathbf{v}_{2}$ $\rightarrow$ $\mathbf{e}_{12}$ $\rightarrow$ $\mathbf{v}_{1}$
(d) Traverse crossing of edges: $\mathbf{v}_{1}$ $\rightarrow$ $\mathbf{e}_{15}$ $\rightarrow$ $\mathbf{v}_{5}$ $\rightarrow$ $\mathbf{e}_{25}$ $\rightarrow$ $\mathbf{v}_{2}$ $\rightarrow$ $\mathbf{e}_{25}$ $\rightarrow$ $\mathbf{v}_{5}$ $\rightarrow$ $\mathbf{e}_{35}$ $\rightarrow$ $\mathbf{v}_{3}$ $\rightarrow$ $\mathbf{e}_{35}$ $\rightarrow$ $\mathbf{v}_{5}$ $\rightarrow$ $\mathbf{e}_{45}$ $\rightarrow$ $\mathbf{v}_{4}$ $\rightarrow$ $\mathbf{e}_{45}$ $\rightarrow$ $\mathbf{v}_{5}$ $\rightarrow$ $\mathbf{e}_{15}$ $\rightarrow$ $\mathbf{v}_{1}$
(e) Traverse closed loop with one vertex: $\mathbf{v}_{1}$ $\rightarrow$ $\mathbf{e}_{11}$ $\rightarrow$ $\mathbf{v}_{1}$
(f) Traverse closed loop with multiple vertices: $\mathbf{v}_{1}$ $\rightarrow$ $\mathbf{e}_{12}$ $\rightarrow$ $\mathbf{v}_{2}$ $\rightarrow$ $\mathbf{e}_{21}$ $\rightarrow$ $\mathbf{v}_{1}$
Figure 15: Exemplary edge graphs to demonstrate the traversal order. Starting from an arbitrary edge in an arbitrary direction, the traversal continues always taking right turns until the loop is closed.
@@
The patch Laplacian $\mathbf{f}(\mathbf{x})$ defines the color gradient inside a patch, which is controlled by evaluating and adding the Laplacian $\mathbf{f}_{\mathbf{p}(t)}(\mathbf{x})$ of Poisson curves $\mathbf{p}(t)\in\mathcal{P}$ and by sampling the color derivatives of gradient meshes, cf. Section 2.2.
$\mathbf{f}(\mathbf{x})=\underbrace{\sum_{\mathbf{p}(t)\in\mathcal{P}}\mathbf{f}_{\mathbf{p}(t)}(\mathbf{x})}_{\textrm{Poisson curves}}+\underbrace{\sum_{\mathbf{c}_{i}(\mathbf{u})\in\mathcal{G}}\lambda_{i}\cdot\Delta_{\mathbf{x}}\mathbf{c}_{i}(\mathbf{u}(\mathbf{x}))}_{\textrm{Gradient meshes}}.$ (12)
If a patch boundary curve of a patch is part of a gradient mesh interior, we add the gradient mesh to a list. Each Laplacian $\Delta_{\mathbf{x}}\mathbf{c}(\mathbf{u}(\mathbf{x}))$ can be evaluated analytically as explained in Appendix A (including the coordinate transformation $\mathbf{u}(\mathbf{x})$ that takes image coordinates to a $(u,v)$ coordinate in the gradient mesh). When a patch is formed from one gradient mesh, then the diffused color gradient inside the patch matches exactly the color gradient of the interpolated gradient mesh. In the event that multiple gradient meshes are overlapping at the pixel coordinate $\mathbf{x}$, we use a weighted sum of the gradient mesh Laplacians to compute the source term $\mathbf{f}(\mathbf{x})$. We offer four choices for how the weights $\lambda_{i}$ are chosen. In the following, let $n$ be the number of gradient meshes, i.e., $i\in\{1,\dots,n\}$.
  1. 1.
    All weights are $\lambda_{i}=0$, i.e., the Laplacians are ignored.
  2. 2.
    All weights are $\lambda_{i}=1$, i.e., all Laplacians are summed.
  3. 3.
    All weights are $\lambda_{i}=\frac{1}{n}$, i.e., the Laplacian’s are averaged.
  4. 4.
    Only one Laplacian is chosen with $\lambda_{1}=1$, and the others are set to zero $\lambda_{2}=\dots=\lambda_{n}=0$.
Options 1) – 3) are order-independent, while option 4) requires the specification of an order. We assume that the gradient meshes $\mathcal{G}$ are given in a particular order by the user, i.e., we choose the gradient mesh that is sitting ’on top’ of all others. For this, the user interface contains a list of all gradient meshes, which can be reordered by the user. Fig. 16 gives examples.
Zero
Sum
Average
First a)
First b)
Figure 16: Different options for handling overlapping gradient meshes. ’Zero’ sets the Laplacian to zero, which gives the smoothest possible interpolation of the boundary conditions. ’Sum’ sums up the Laplacians, which can lead to high values when many gradient meshes overlap. ’Average’ computes the average of the Laplacians, which scales better to larger numbers of gradient meshes. ’First’ only takes one of the Laplacians, for which we show both options: choosing the smooth square in a) and the striped square in b).